OpenStudy (anonymous):

Solve P = 2l + 2w for w in terms of P and l.

OpenStudy (ivettef365):

to solve for w, the way you isolate as follows: P = 2l + 2w (first move the 2l by changing sign as follows) P-2l = 2w

OpenStudy (ivettef365):

now you have P-2l = 2w (now divide by 2 on both sides) then you have ---- --- 2 2

OpenStudy (ivettef365):

so P-2l w = ----- 2